I don't understand this question.
A computer is infected with the Sasser viruse. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another 5 minutes, etc. How long until 100 million computers are infected?

Answer by scott8148(6628)

every 5 minutes, the number of infected computers is MULTIPLIED by 20
__ 1...20...400...8000...etc __ how long until 100 million?

let t="time in minutes" __ 100M=1(20)^(t/5)

taking log __ log(100M)=(t/5)(log(20))

dividing by log(20) __ log(100M)/log(20)=t/5

multiplying by 5 __ 5[log(100M)/log(20)]=t